If f(2) 4 – x2 3 + x2 find: f(x) =
Answer #1

a) f(x)= 4-x? 3 ta? So, df (3+2²) (2x)-(4-2²) (22) da ( 3+x) 2 -62 -22 3-8x + 2x² . f(x) -14% (3+2?)? f(x)= 3 + 2)2 Kindly re